1. Imagine that 50 years ago, your aunt made an investment of $10 into an account that earned 12% each year. What is the effect of compounding? (Round to 2 decimal places, do not include dollar signs in your response.)

2. When first introduced in 1985, Nike Air Jordan’s sold for $65. Today (July 2, 2018) on stockx.com, a pair sold for $3023. What is the annual rate of growth of the price of the Air Jordan’s? (Note – assume t=33 years.) Include only numbers in your answer. Convert to percent. Round to 2 decimal places.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.A=$10*(1.12)^50

=2890.02(Approx).

2.

3023=65*(1+r/100)^33

(3023/65)^(1/33)=(1+r/100)

(1+r/100)=1.1234

r=1.1234-1

=12.34(Approx).